15. [lpt] A 59.6 kg bungee jumper jumps from a bridge. She is tied to a 12.4 m long bungee cord and falls a total of 34.2 m. Calculate the spring constant k of the bungee cord. Answer: Submit All Answers
Answer #1

c longahon o chd = 342-12.4 ne 18.6x9-65 26.88N/m. Spring consfauf C = 26* Pué N/m